ABCD is a square and P, Q are the mid points of BC, CD respectively. If AP = a and AQ = b, find in terms of a and b, the directed line segments 

I. AB

II. AD

III. BD 

IV. AC


1
Expert's answer
2021-01-25T03:17:14-0500

"Let\:ABCD\:be\:a\:square."

"Let\:equal\:sides\:of\:square=x"

"Given\:that\:P\:and\:Q\:are\:middle\:points\:of\:BC\:and\:CD\:respectively"

"BC=CD=x"

"BP=\frac{x}{2}\:and\:QD\:=\frac{x}{2}"

"Given\:that\:AP=a\:and\:AQ=b\:"

"Now\:in\:right\:angled\:triangle\:ABP"

"By\:pythagorous\:theorem"

"\left(AP\right)^2=\left(AB\right)^2+\left(BP\right)^2"

"a^2=x^2+\left(\frac{x}{2}\right)^2"

"a^2=\frac{5x^2}{4}"

"Now\:in\:right\:angled\:triangle\:ADQ"

"By\:pythagorous\:theorem"

"\left(AQ\right)^2=\left(AD\right)^2+\left(DQ\right)^2"

"b^2=x^2+\frac{x^2}{4}"

"b^2=\frac{5x^2}{4}"


"a^2=b^2=\frac{5x^2}{4}\:or\:a=b=\frac{\sqrt{5x}}{2}"

"\:x=\frac{2a}{\sqrt{5}}\:or\:x=\frac{2b}{\sqrt{5}}"

"I.\:AB=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}"

"II.\:AD=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}"

"III.\:For\:BD"

"\left(BD\right)^2=\left(BC\right)^2+\left(CD\right)^2"

"=x^2+x^2"

"=2x^2"

"=2\times\:\frac{4a^2}{5}"

"BD=\sqrt{2}\times\frac{2a}{\sqrt{5}}\:or\:\sqrt{2}\times\frac{2b}{\sqrt{5}}"

"IV.\:AC"

"For\:AC"

"\left(AC\right)^2=x^2+x^2=2x^2"

"\:AC=\frac{2\sqrt{2a}}{\sqrt{5}}\:or\:\frac{2\sqrt{2b}}{\sqrt{5}}"